A 3% solution of sulfuric acid was mixed with an 18% solution of sulfuric acid to produce an 8% solution. How much 3% and how much 18% solution were used to produce 15L of 8% solution?
Answer by checkley79(3341)

.18x+.03(15-x)=.08*15
.18x+.45-.03x=1.2
.15x=1.2-.45
.15x=.75
x=.75/.15
x=5 L of 18% solution is used.
15-5=10 L of 3% solution is used.
Proof:
.18*5+.03*10=.08*15
.90+.30=1.20
1.20=1.20